A solid cylinder of height h cm and diameter 14 cm is melted and recast into two cones of the same height in the proportion of 3 :4 of volume. Find the percentage change in the surface area excluding the curved surfaces

Grade:10

1 Answers

Don Boii
6 years ago
2+2=4-1=3 quick maffssigsigdohdohdoufpufoufouc9te95e96w9totsigdogdob jpvljcojfohdohdoydiydiysitwotwotsotsogdogdohfohfohdohdohdohdogsogclhxohdysogsohditwitkgdurwigxjfwurdkgxoueurw